a right cylindrical container has a radius of 12 centimeters and the height of 15 centimeters a right conical containers base has a diameter of 24 centimeters if the conical container hase the same volume as the cylindrical container what it's the height of the conical container
Answer by amarjeeth123(569)

Radius of cylinder=12cm
Height of cylinder=15cm
Volume of cylinder=π*r^2*h=π*(12)^2*(15)=2160π cubic cm.
Diameter of cone=24 cm
Radius of cone=12 cm
Let the height of the cone be h.
Volume of cone=1/3*π*r^2*h=1/3*π*(12)^2*h
Volume of cylinder=Volume of cone
π*(12)^2*(15)=1/3*π*(12)^2*h
Simplifying we get h=45 cm.
Height of the cone is 45 cm.